# Extended Emission Line Region in a Post-Starburst Galaxy Hosting TDE AT2019qiz and QPE

### v2

@@ Abstract
We present a comprehensive analysis of the extended emission line region (EELR) in the host galaxy of the tidal disruption event (TDE) AT2019qiz, utilizing VLT/MUSE integral-field spectroscopy. The high spatial-resolution data reveal a bi-conical emission structure approximately in scale within the galactic center, characterized by a prominent [O iii] line in the nucleus and significant [N ii] line emission extending into the EELR. Spectral analysis of the EELR indicates line ratios consistent with Seyfert ionization in the center and LINER-type ionization in the outer diffuse region, suggesting ionization from galactic nuclear activity. The required ionizing luminosity, estimated from the H and H luminosities based on the photoionization and recombination balance assumption, is for all spaxels classified as AGN, and for spaxels in the central Seyfert region. However, the current bolometric luminosity of the nucleus , estimated from quiescent-state soft X-ray observations, is insufficient to ionize the entire EELR, implying a recently faded AGN or a delayed response to historical activity. Stellar population analysis reveals a post-starburst characteristic in the EELR, and the gas kinematics show disturbances and non-circular components compared to the stellar kinematics. Notably, the recent detection of quasi-periodic eruptions (QPEs) in the X-ray light curve of AT2019qiz confirms the TDE-QPE association. Our findings provide direct evidence for an AGN-like EELR in the host galaxy of the nearest TDE with QPE detection, offering new insights into the complex interplay between TDEs, QPEs, AGN activity, and host galaxy evolution.
UTF8gbsn
@@ 1 INTRODUCTION
Tidal disruption event (TDEs) and quasi-periodic eruptions (QPEs) are two distinct transient phenomena observed in galactic nuclei in recent years. TDEs are well known phenomena which occur when stars are tidally disrupted by supermassive black holes (SMBHs) (Rees, 1988; Gezari, 2021), producing characteristic light curves across multiple wavelengths. QPEs, on the other hand, are characterized by nearly periodic X-ray eruptions with recurrence times of hours to days, whose physical origin remains debated. Current theoretical explanations for QPEs primarily include instabilities originating from the accretion disk itself (e.g. Sniegowska et al., 2020; Raj & Nixon, 2021; Pan et al., 2022, 2023), mass transfer from a partially disrupted star or white dwarf via Roche-lobe overflow (e.g. King, 2022; Krolik & Linial, 2022; Wang et al., 2022; Metzger et al., 2022; Linial & Sari, 2023), and periodic impacts of a single stellar mass object with the accretion disk of the SMBH, such as a main sequence star or a stellar mass black hole (e.g. Lu & Quataert, 2023; Linial & Metzger, 2023; Franchini et al., 2023; Tagawa & Haiman, 2023; Zhou et al., 2024a, c, b). All theoretical models require the presence of an accretion disk, which can form either through the direct accretion of gas by the SMBHs or as a consequence of the tidal disruption of stars during a TDE. Recent observations suggest a potential connection between TDEs and QPEs. For instance, TDE-like declining light curves have been observed in QPE events (GSN069; Shu et al. (2018); Miniutti et al. (2023), eRO-QPE3; Arcodia et al. (2024)), and QPE-like eruptions have been detected in two TDEs (XMMSL1 J024916.6-04124; Chakraborty et al. (2021), AT 2019vcb; Quintin et al. (2023)). The most important progress recently was that Nicholl et al. (2024) observed nine quasi-periodic X-ray eruptions from 29 February 2024 to 14 March 2024 in the host galaxy of TDE AT2019qiz, which is one of the nearest TDE event discovered on 2019-9-19 UT and reach its peak luminosity at 2019-10-08 UT (Nicholl et al., 2020). This is the first confirmed repeating QPE in a spectroscopically confirmed TDE.
@@
Recent studies about the host galaxy properties of the TDEs (e.g. French et al., 2020) and the QPEs (e.g. Wevers et al., 2022) have revealed intriguing similarities between them (Wevers et al., 2024). Both types of events predominantly occur in post-starburst or E+A galaxies (e.g. French et al., 2016, 2017) characterized by high stellar surface mass densities, Sérsic indices, and high bulge-to-total light (B/T) ratios (Law-Smith et al., 2017; Graur et al., 2018; Gilbert et al., 2024). Intriguingly, extended emission line regions (EELRs), also named as extended narrow line regions (ENLRs), consisting of ionized gas spanning scales of several to ten kpc, have been detected in several host galaxies of both TDEs and QPEs. To date, three TDEs (ASASSN–14li; Prieto et al. (2016), AT2019azh; French et al. (2023), and iPTF–16fnl; Wevers & French (2024)) have been reported to exhibit EELRs in their host post-starburst galaxy. Among the ten known QPEs and candidates (Sun et al., 2013; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022, 2024; Chakraborty et al., 2021; Evans et al., 2023; Guolo et al., 2024; Nicholl et al., 2024), three of them (GSN 069; Patra et al. (2024), RXJ1301, and eRO-QPE2; Wevers et al. (2024)) have been detected with EELRs using Hubble Space Telescope (HST) narrow-band images and the MUSE integral-field spectrograph. The fractional incidence of EELRs in QPE and TDE host galaxies appears higher than in the general galaxy population, despite the limited sample size (Wevers & French, 2024; Wevers et al., 2024). All these EELRs require ionization from a non-stellar continuum source, such as radiation from AGN accretion disks.
@@
However, the nuclear luminosities currently observed in these sources are too low to maintain the necessary ionization (Wevers & French, 2024; Wevers et al., 2024), pointing toward a recently faded AGN scenario (Keel et al., 2017; French et al., 2023). Furthermore, some theoretical models require the presence of an accretion disk prior to the onset of QPEs. For example, the leading EMRI+disk model (Linial & Metzger, 2023; Franchini et al., 2023) suggests that an extreme mass ratio inspiral (EMRI), in which a stellar-mass object orbits a SMBH, collides with an accretion disk formed from a TDE twice per orbit. Detailed orbital analyses show that a large fraction of QPE EMRIs exhibit low orbital eccentricity, supporting the hypothesis that they were born in or were captured by a past AGN disk (Zhou et al., 2024a, c, b), named wet EMRI formation channel (Sigl et al., 2007; Levin, 2007; Pan & Yang, 2021; Pan et al., 2021). In short, these findings suggest a potential physical connection among host galaxy properties, past nuclear activity, TDEs, and the occurrence of QPEs. The burst of star formation together with the galactic nuclear activities and the elevated TDE rate produces UV photons from accretion disks, leading to the ionization of surrounding gas.
@@ 2 MUSE Data and Analysis
We obtained the integral-field spectrograph VLT MUSE (Bacon et al., 2010) Phase3 datacube from the ESO archive, which was observed on 2021-02-08 (Program ID:105.20GS, PI: Sam Kim), 481 rest-frame days after the TDE reach its peak luminosity, with an effective exposure time of 2999 s. The spectral coverage ranges from 470 to 935 nm with a spectral resolution of R = 3014. The observation was conducted in WFM-AO-N mode, and the sodium light contaminated region between 582 and 597 nm was excluded. The spatial resolution is /spaxel, and the DIMM seeing is . We performed multiple analyses on the datacube. First, we generated emission line maps by integrating the continuum-subtracted spectra around each emission line, as shown in Section 2.1. We then conducted detailed spectral fitting on both individual regions of interest and the entire datacube after 3×3 spatial binning to improve the signal-to-noise ratio (SNR), which will be presented in Section 2.2.
@@ 2.1 Datacube Quicklook: An Extended Emission Line Region
We first performed a quick inspection of the datacube. For each emission line, we obtained line flux distribution maps by integrating the continuum-subtracted emission line profiles, where the continuum was estimated through linear interpolation of the adjacent spectral regions. The [O iii]5007, H and [N ii]6584 maps are presented in the Appendix Figure A.1. This straightforward method helps avoid potential artifacts from poor Gaussian line fitting. The maps reveal a prominent extended emission line region at the center of the host galaxy, particularly evident in the [N ii] emission. In Figure 1, we cropped the datacube of the galaxy region, and made a composite RGB color image to illustrate the distribution of the extended emission line region relative to the host galaxy. This color image was created by convolving the datacube with Bessel R, V, and B filter transmission curves from the SVO Filter Profile Service (Rodrigo & Solano, 2020). The integrated [N ii] flux map is overlaid on the color image of the galaxy. The extended emission line region appears to be distributed within the galactic bar at the center, exhibiting a bi-conical structure extending approximately , the boundary is estimated just by visual inspection.
@@ 2.2 Detailed Spectral Fitting and Datacube Analysis
Prior to fitting, all spectra were corrected for Galactic extinction (; Fitzpatrick, 1999; Schlafly & Finkbeiner, 2011). We utilized the pPXF code (Cappellari, 2017, 2023) to perform simultaneous fitting of the stellar population continuum and emission lines across the spectral range from 4700 Å to 7400 Å. The continuum spectra were fitted using E-MILES stellar library (Vazdekis et al., 2010, 2015, 2016) and multiplied by the attenuation curve of Calzetti et al. (2000) to account for the dust attenuation of the host galaxy. Sky line contamination was evident in the spectra, so we constructed a sky spectrum template from the sky background of the datacube and included an 8th-order additive polynomial in the fitting. Narrow emission lines (H, H, [N ii], [S ii], [O iii]) were modeled using line templates with an average instrument FWHM of 2.65 Å, convolved with a Gaussian line-of-sight velocity distribution (LOSVD) function. For spectra within the white dashed circle in Figure 1, two broad Balmer line components for each H and H lines were included in the fitting. Through this fitting procedure, we obtained measurements of emission line fluxes, stellar population properties including age and metallicity, as well as the kinematics of both gas and stars. The spectra for each region and the fitting results are presented in Figure 2 for the entire EELR and Figure A.2 for Region 1,2,3 and center. It should be noted that the narrow H lines cannot be seen in the original spectral data, as there are only absorption lines. However, they can be revealed after subtracting the modeled stellar continuum. We believe this may overestimate the flux of the H lines, so the [O iii]/H ratio in the BPT diagram may be underestimated. Similarly, the H emission lines exhibit lower flux compared to [N ii] in the original data. After correcting for the absorption feature, the residual H emission will show enhanced flux levels than before. We also performed measurements on the entire datacube. To improve fitting efficiency, we cropped the sky background region without useful signal, focusing only on the region of the galaxy, and applied 3 3 spaxel binning to the original datacube to enhance the signal-to-noise ratio, which degrade the spatial resolution to /spaxel. We then conducted batch measurements using pPXF with the same settings as mentioned above. The broad line components in the central spaxels were also considered.
@@ 3.1 BPT Diagrams and Stellar Population of the EELR
Additionally, the stellar population analysis reveals a star formation event approximately 1 Gyr ago, as indicated by the fitting results from the entire EELR region, as well as local Regions 1, 2, 3, and the central region. Figure A.3 illustrates the stellar population result for the entire EELR region. The pPXF fitting enables us to derive weights for various E-MILES stellar population spectral templates (Vazdekis et al., 2016). Consequently, we can plot weight maps for ages and metallicities, and display weights for different ages, which can be interpreted as the star formation history. It is evident that after the initial star formation 10 Gyr ago, another star formation event occurred 1 Gyr ago, after which star formation was quenched. This pattern is characteristic of post-starburst galaxies. As mentioned in the introduction, TDEs and QPEs are more likely to occur in post-starburst galaxies, a finding corroborated by our analysis. However, the spectral wavelength range of the MUSE spectrum does not include or for our galaxy. Future spectral observations with broader wavelength coverage are necessary for more reliable stellar population synthesis.
@@ 3.2 Emission Line Flux and BPT Classification Map
It is worth noting that the ionization mechanism of spectra classified as LINER remains controversial. In addition to ionization by low-luminosity AGNs (Ho, 2008), it may also originate from ultraviolet photons produced by evolved stellar populations (such as post-AGB stars) (Yan & Blanton, 2012; Singh et al., 2013; Belfiore et al., 2016, e.g.) or shocks front passing through the ISM (Heckman, 1980; Dopita et al., 2015). TDEs with an elevated event rate have also been considered as an alternative ionizing source of EELRs in TDE host galaxies (Wevers & French, 2024). A recent study (Mummery et al., 2025) shows that ionizing radiation flux from a TDE disk with a lifetime of yr may power a galactic scale EELR that lasts for yr, and contaminates a fraction of AGN identifications in BPT diagrams. Therefore, the ionization mechanism of the diffuse regions in the outer EELR warrants further modeling and analysis in the future. Specifically, studies combining spatially resolved stellar population types and luminosity measurements together with photoionization models could provide valuable insights.

### v4

# Extended Emission Line Region in a Post-Starburst Galaxy Hosting Tidal Disruption Event AT2019qiz and Quasi-periodic Eruptions

@@ Abstract
We present a comprehensive analysis of the extended emission line region (EELR) in the host galaxy of the tidal disruption event (TDE) AT2019qiz, utilizing Very Large Telescope/MUSE integral-field spectroscopy. The high spatial-resolution data reveal a bi-conical emission structure approximately in scale within the galactic center, characterized by a prominent [O iii] line in the nucleus and significant [N ii] line emission extending into the EELR. Spectral analysis of the EELR indicates line ratios consistent with Seyfert ionization in the center and LINER-type ionization in the outer diffuse region, suggesting ionization from galactic nuclear activity. The required ionizing luminosity, estimated from the H and H luminosities based on the photoionization and recombination balance assumption, is for all spaxels classified as active galactic nucleus (AGN), and for spaxels in the central Seyfert region. However, the current bolometric luminosity of the nucleus , estimated from quiescent-state soft X-ray observations, is insufficient to ionize the entire EELR, implying a recently faded AGN or a delayed response to historical activity. Stellar population analysis reveals a post-starburst characteristic in the EELR, and the gas kinematics show disturbances and non-circular components compared to the stellar kinematics. Notably, the recent detection of quasi-periodic eruptions (QPEs) in the X-ray light curve of AT2019qiz confirms the TDE-QPE association. Our findings provide direct evidence for an AGN-like EELR in the host galaxy of the nearest TDE with QPE detection, offering new insights into the complex interplay between TDEs, QPEs, AGN activity, and host galaxy evolution.
@@ 1 INTRODUCTION
Tidal disruption events (TDEs) and quasi-periodic eruptions (QPEs) are two distinct transient phenomena observed in galactic nuclei in recent years. TDEs are well known phenomena which occur when stars are tidally disrupted by supermassive black holes (SMBHs) (Rees, 1988; Gezari, 2021), producing characteristic light curves across multiple wavelengths. QPEs, on the other hand, are characterized by nearly periodic X-ray eruptions with recurrence times of hours to days, whose physical origin remains debated. Current theoretical explanations for QPEs primarily include instabilities originating from the accretion disk itself (e.g. Sniegowska et al., 2020; Raj & Nixon, 2021; Pan et al., 2022, 2023), mass transfer from a partially disrupted star or white dwarf via Roche-lobe overflow (e.g. King, 2022; Krolik & Linial, 2022; Wang et al., 2022; Metzger et al., 2022; Linial & Sari, 2023), and periodic impacts of a single stellar mass object with the accretion disk of the SMBH, such as a main sequence star or a stellar mass black hole (e.g. Lu & Quataert, 2023; Linial & Metzger, 2023; Franchini et al., 2023; Tagawa & Haiman, 2023; Zhou et al., 2024a, b, 2025). All theoretical models require the presence of an accretion disk, which can form either through the direct accretion of gas by the SMBHs or as a consequence of the tidal disruption of stars during a TDE. Recent observations suggest a potential connection between TDEs and QPEs. For instance, TDE-like declining light curves have been observed in QPE events (GSN069; Shu et al. 2018; Miniutti et al. 2023, eRO-QPE3; Arcodia et al. 2024), and QPE-like eruptions have been detected in two TDEs (XMMSL1 J024916.6-04124; Chakraborty et al. 2021, AT2019vcb; Quintin et al. 2023). The most important progress recently was that Nicholl et al. (2024) observed nine quasi-periodic X-ray eruptions from 29 February 2024 to 14 March 2024 in the host galaxy of TDE AT2019qiz, which is one of the nearest TDE discovered on 2019-9-19 UT and reach its peak luminosity at 2019-10-08 UT (Nicholl et al., 2020). This is the first confirmed repeating QPE in a spectroscopically confirmed TDE. After that, repeated QPEs after TDE have also been reported in AT2022upj (Chakraborty et al., 2025).
@@
Recent studies about the host galaxy properties of the TDEs (e.g. French et al., 2020) and the QPEs (e.g. Wevers et al., 2022) have revealed intriguing similarities between them (Wevers et al., 2024). Both types of events predominantly occur in post-starburst or E+A galaxies (e.g. French et al., 2016, 2017) characterized by high stellar surface mass densities, Sérsic indices, and high bulge-to-total light (B/T) ratios (Law-Smith et al., 2017; Graur et al., 2018; Gilbert et al., 2024). Intriguingly, extended emission line regions (EELRs) (e.g. Lintott et al., 2009; Keel et al., 2012, 2015), also named as extended narrow line regions (ENLRs), consisting of ionized gas spanning scales of several to ten kpc, have been detected in several host galaxies of both TDEs and QPEs. To date, three TDEs (ASASSN–14li; Prieto et al. (2016), AT2019azh; French et al. (2023), and iPTF–16fnl; Wevers & French (2024)) have been reported to exhibit EELRs in their host post-starburst galaxy. Among the about 12 known QPEs and candidates (Sun et al., 2013; Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021, 2022, 2024; Chakraborty et al., 2021; Evans et al., 2023; Guolo et al., 2024; Nicholl et al., 2024; Hernández-Garc´ıa et al., 2025; Chakraborty et al., 2025), three of them (GSN 069; Patra et al. (2024), RXJ1301, and eRO-QPE2; Wevers et al. (2024)) have been detected with EELRs using Hubble Space Telescope (HST) narrow-band images and the MUSE integral-field spectrograph. The fractional incidence of EELRs in QPE and TDE host galaxies appears higher than in the general galaxy population (Wevers & French, 2024; Wevers et al., 2024), despite the limited sample size and the fact that MUSE observations provide deeper detection thresholds than narrowband-imaging surveys (e.g. Keel et al., 2024), making a comparison less direct.
@@
All the EELRs in QPE and TDE host galaxies require ionization from a non-stellar continuum source, such as radiation from active galactic nucleus (AGN) accretion disks. However, the nuclear luminosities currently observed in these sources are too low to maintain the necessary ionization (Wevers & French, 2024; Wevers et al., 2024), pointing toward a recently faded AGN scenario (Keel et al., 2017; French et al., 2023). Furthermore, some theoretical models require the presence of an accretion disk prior to the onset of QPEs. For example, the leading EMRI+disk model (Linial & Metzger, 2023; Franchini et al., 2023) suggests that an extreme mass ratio inspiral (EMRI), in which a stellar-mass object orbits a SMBH, collides with an accretion disk formed from a TDE twice per orbit. Detailed orbital analyses show that a large fraction of QPE EMRIs exhibit low orbital eccentricity, supporting the hypothesis that they were born in or were captured by a past AGN disk (Zhou et al., 2024a, b, 2025), named wet EMRI formation channel (Sigl et al., 2007; Levin, 2007; Pan & Yang, 2021; Pan et al., 2021). In short, these findings suggest a potential physical connection among host galaxy properties, past nuclear activity, TDEs, and the occurrence of QPEs. The burst of star formation together with the galactic nuclear activities and the elevated TDE rate produces UV photons from accretion disks, leading to the ionization of surrounding gas.
@@ 2 MUSE Data and Analysis
We obtained the integral-field spectrograph VLT MUSE (Bacon et al., 2010) Phase3 datacube from the ESO archive, which was observed on 2021-02-08 (Program ID:105.20GS, PI: Sam Kim), 481 rest-frame days after the TDE reach its peak luminosity, with an effective exposure time of 2999 s. The spectral coverage ranges from 470 to 935 nm with a spectral resolution of R = 3014. The observation was conducted in WFM-AO-N mode, and the sodium light contaminated region between 582 and 597 nm was excluded. The spatial resolution is , and the DIMM seeing is . We performed multiple analyses on the datacube. First, we generated emission line maps by integrating the continuum-subtracted spectra around each emission line, as shown in Section 2.1. We then conducted detailed spectral fitting on both individual regions of interest and the entire datacube after 3×3 spatial binning to improve the signal-to-noise ratio (SNR), which will be presented in Section 2.2.
@@ 2.1 Datacube Quicklook: An Extended Emission Line Region
We first performed a quick inspection of the datacube. For each emission line, we obtained line flux distribution maps by integrating the continuum-subtracted emission line profiles, where the continuum was estimated through linear interpolation of the adjacent spectral regions. The [O iii]5007, H and [N ii]6584 maps are presented in the Appendix Figure A.1. This straightforward method helps avoid potential artifacts from poor Gaussian line fitting. The maps reveal a prominent EELR at the center of the host galaxy, particularly evident in the [N ii] emission. In Figure 1, we cropped the datacube of the galaxy region, and made a composite RGB color image to illustrate the distribution of the EELR relative to the host galaxy. This color image was created by convolving the datacube with Bessel R, V, and B filter transmission curves from the SVO Filter Profile Service (Rodrigo & Solano, 2020). The integrated [N ii] flux map is overlaid on the color image of the galaxy. The EELR appears to be distributed within the galactic bar at the center, exhibiting a bi-conical structure extending approximately , the boundary is estimated just by visual inspection.
@@ 2.2 Detailed Spectral Fitting and Datacube Analysis
Prior to fitting, all spectra were corrected for Galactic extinction (; Fitzpatrick, 1999; Schlafly & Finkbeiner, 2011). We utilized the pPXF code (Cappellari, 2017, 2023) to perform simultaneous fitting of the stellar population continuum and emission lines across the spectra range from 4700 Å to 7400 Å. The continuum spectra were fitted using E-MILES stellar library (Vazdekis et al., 2010, 2015, 2016) and multiplied by the attenuation curve of Calzetti et al. (2000) to account for the dust attenuation of the host galaxy. Sky line contamination was evident in the spectra, so we constructed a sky spectrum template from the sky background of the datacube and included an 8th-order additive polynomial in the fitting. Narrow emission lines (H, H, [N ii], [S ii], [O iii]) were modeled using line templates with an average instrument FWHM of 2.65 Å, convolved with a Gaussian line-of-sight velocity distribution (LOSVD) function. For spectra within the white dashed circle in Figure 1, two broad Balmer line components for each H and H lines were included in the fitting. Through this fitting procedure, we obtained measurements of emission line fluxes, stellar population properties including age and metallicity, as well as the kinematics of both gas and stars.
The spectra for each region and the fitting results are presented in Figure 2 for the entire EELR and Figure A.2 for Region 1,2,3 and center. It should be noted that the narrow H lines cannot be seen in the original spectral data, as there are only absorption lines. However, they can be revealed after subtracting the modeled stellar continuum. We believe this may overestimate the flux of the H lines, so the [O iii]/H ratio in the BPT diagram may be underestimated. Similarly, the H emission lines exhibit lower flux compared to [N ii] in the original data. After correcting for the absorption feature, the residual H emission will show enhanced flux levels than before. The He ii4686 line, which is also used as an AGN indicator (e.g. Shirazi & Brinchmann, 2012; Tozzi et al., 2023) and can be quite strong in some EELRs (Keel et al., 2012), lies near the edge of our wavelength range when redshifted (4756 Å). We detect only a broad He ii4686 component in the galactic centre, which interpreted as an outflow from AT2019qiz (Nicholl et al., 2020), while the EELR shows no significant He ii4686 emission. The position of the line close to the wavelength limit makes our stellar population continuum fitting poor, and also prevents an accurate measurement of the He ii4686 flux. Therefore, we omit the analysis of He ii4686 in this paper. We also performed measurements on the entire datacube. To improve fitting efficiency, we cropped the sky background region without useful signal, focusing only on the region containing the galaxy, and applied 3 3 spaxel binning to the original datacube to enhance the signal-to-noise ratio, which degrade the spatial resolution to . We then conducted batch measurements using pPXF with the same settings as mentioned above. The broad line components in the central spaxels were also considered.
@@ 3.1 BPT Diagrams and Stellar Population of the EELR
Additionally, the stellar population analysis reveals a star formation event approximately 1 Gyr ago, as indicated by the fitting results from the entire EELR region, as well as local Regions 1, 2, 3, and the central region. Figure A.3 illustrates the stellar population result for the entire EELR region. The pPXF fitting enables us to derive weights for various E-MILES stellar population spectral templates (Vazdekis et al., 2016). Consequently, we can plot weight maps for ages and metallicities, and display weights for different ages, which can be interpreted as the star formation history. It is evident that after the initial star formation 10 Gyr ago, another star formation event occurred 1 Gyr ago, after which star formation was quenched. The different weights of the stellar spectra correspond to different mass-to-light ratios. Through this analysis, we estimate that this star formation event approximately 1 Gyr ago formed stars accounting for about 10% of the galaxy’s total stellar mass. This pattern is characteristic of post-starburst galaxies (Bergvall et al., 2016). We also found that the host of AT2019qiz is located in the “green valley” region, approaching the “red sequence” in the galaxy color-magnitude diagram (Bell et al., 2004). This classification is based on g and r band data (after corrections for Galactic extinction) from the DESI Legacy Survey (Dey et al., 2019), which is consistent with the classification in Hammerstein et al. (2021). As mentioned in the introduction, TDEs and QPEs are more likely to occur in post-starburst galaxies, a finding corroborated by our analysis. However, the spectral wavelength range of the MUSE spectrum does not include or for our galaxy. Future spectral observations with broader wavelength coverage are necessary for more reliable stellar population synthesis.
@@ 3.2 Emission Line Flux and BPT Classification Map
It is worth noting that the ionization mechanism of spectra classified as LINER remains controversial. In addition to ionization by low-luminosity AGNs (Ho, 2008), it may also originate from ultraviolet photons produced by evolved stellar populations (such as post-AGB stars) (e.g. Yan & Blanton, 2012; Singh et al., 2013; Belfiore et al., 2016) or shocks front passing through the ISM (Heckman, 1980; Dopita et al., 2015). TDEs with an elevated event rate have also been considered as an alternative ionizing source of EELRs in TDE host galaxies (Wevers & French, 2024). A recent study (Mummery et al., 2025) shows that ionizing radiation flux from a TDE disk with a lifetime of yr may power a galactic scale EELR that lasts for yr, and contaminates a fraction of AGN identifications in BPT diagrams. Therefore, the ionization mechanism of the diffuse regions in the outer EELR warrants further modeling and analysis in the future. Specifically, studies combining spatially resolved stellar population types and luminosity measurements together with photoionization models could provide valuable insights.